OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y=3x+6 and passes trhough 4,7

OpenStudy (anonymous):

The answer is y=3x-5. Using point-slope formula: y-y1=m(x-x1) m=3 coordinates: (4,7) y-7=3(x-4) y-7=3x-12 y=3x-12+7 y=3x-5